# Wearable Accelerometer Foundation Models for Health via Knowledge Distillation

### v1

@@ Abstract
Modern wearable devices can conveniently and continuously record various biosignals in the many different environments of daily living, ultimately enabling a rich view of individual health. However, not all biosignals are the same: high-fidelity measurements, such as photoplethysmography (PPG), contain more physiological information, but require optical sensors with a high power footprint. In a resource-constrained setting, such biosignals may be unavailable. Alternatively, a lower-fidelity biosignal, such as accelerometry that captures minute cardiovascular information during low-motion periods, has a significantly smaller power footprint and is available in almost any wearable device. Here, we demonstrate that we can distill representational knowledge across biosignals, i.e., from PPG to accelerometry, using 20 million minutes of unlabeled data, collected from $\sim$172K participants in the Apple Heart and Movement Study under informed consent. We first pre-train PPG encoders via self-supervised learning, and then distill their representational knowledge to accelerometry encoders. We demonstrate strong cross-modal alignment on unseen data, e.g., $99.2\%$ top-1 accuracy for retrieving PPG embeddings from accelerometry embeddings. We show that distilled accelerometry encoders have significantly more informative representations compared to self-supervised or supervised encoders trained directly on accelerometry data, observed by at least 23%-49% improved performance for predicting heart rate and heart rate variability. We also show that distilled accelerometry encoders are readily predictive of a wide array of downstream health targets, i.e., they are generalist foundation models. We believe accelerometry foundation models for health may unlock new opportunities for developing digital biomarkers from any wearable device, and help individuals track their health more frequently and conveniently.
@@ 1 Introduction
The recent growth of wearable devices has empowered individuals to track their health more conveniently and frequently. Wearable devices can record various biosignals and provide individuals with a continuous view of their health metrics in daily environments, a view which is hard to achieve in clinical settings. While this unlocks new exciting opportunities for wearable devices, challenges remain. High-fidelity biosignals that have substantial physiological information content tend to require specific hardware and sensors for recording, therefore making them unavailable in some wearable devices. Even when available, they may be difficult to collect frequently due to power constraints of small wearable devices, or they may require active user engagement. An example of such is photoplethysmography (PPG), which measures volumetric changes in arterial blood flow and contains diverse physiological information but requires power-consuming optical sensors for recording. On the other hand, lower-fidelity biosignals, such as accelerometry, are available on almost all wearable devices and has a significantly small power footprint, which makes them ideal for enabling digital biomarkers efficiently and on a variety of wearable devices. While accelerometry during gross motion is widely used for activity recognition and fitness, our focus is on accelerometry during low motion. During low-motion periods, accelerometry is capable of detecting minute physiological movements on the body as a result of heart function, known as ballistocardiogram (Inan et al., 2015; Kim et al., 2016), but its relationship to diverse health targets is not well-understood, is susceptible to motion artifacts and challenging to model. Examples of how less-structured and noisier accelerometry is compared to PPG can be found in Figure 1, top right. As a result, despite its ubiquity, accelerometry has been under-explored for health diagnosis and developing digital biomarkers for health on wearable devices. See Section 2 for how our work relates to the existing work.
@@
In this paper, using unlabeled sensor data collected under informed consent from the large longitudinal Apple Heart and Movement Study (AHMS) (MacRae, 2021; Truslow et al., 2024), we demonstrate that we can train generalist accelerometry foundation models via cross-modal representational knowledge distillation from PPG. Our contributions are: 1) Accelerometry foundation models for health: We train accelerometry foundation models for health via representational knowledge distillation from PPG using a large-scale dataset with 20M minutes of multi-modal sensor data from $\sim$172K participants. We show that a single accelerometry encoder is predictive of a wide array of downstream health targets with a high accuracy. 2) Fully unsupervised representational knowledge distillation framework: While multi-modal modeling and cross-modal reconstruction of biosignals have been done before, we combine and adopt techniques inspired by uni-modal and multi-modal pre-training frameworks from other domains of deep learning to create a fully unsupervised distillation framework for biosignal time-series, which is particularly crucial for health applications where labeled data is limited. 3) Studying representational alignment: We study representational alignment of PPG and accelerometry encoders, by doing retrieval analysis. 4) Studying representational information: We study the representational power of accelerometry encoders for critical health targets such as heart rate and heart rate variability, across different available labeled data regimes, as well as for demographic variables and 46 health targets including health conditions, use of medications and lifestyle habits from AHMS survey questions. In addition to these contributions, to support the robustness of the findings, we perform ablation studies to evaluate the efficacy of the encoder architecture (Transformer and EfficientNet), the pre-training strategy of the PPG teacher model (masked autoencoding and contrastive learning), the augmentations, and other training choices.
@@ 2 Related work
Uni-modal and multi-modal representation learning: Unsupervised representation learning techniques, e.g., self-supervised learning, have been proven successful in training generalist models, also known as foundation models, without requiring any explicit labels during training in various domains of deep learning such as natural language processing (Devlin et al., 2019; OpenAI, 2023), computer vision (Chen et al., 2020; 2021; He et al., 2021; Oquab et al., 2023), speech recognition (Baevski et al., 2020; 2022), and health (Cheng et al., 2020; Kostas et al., 2021; Sarkar & Etemad, 2022; Mohsenvand et al., 2020; Gopal et al., 2021; Kiyasseh et al., 2021; Mehari & Strodthoff, 2022; Wu et al., 2020; Spathis et al., 2021; Tang et al., 2021; Yuan et al., 2023; Lai et al., 2023; Abbaspourazad et al., 2024; Liu et al., 2024; Narayanswamy et al., 2024). While most of these works have been primarily on training uni-modal foundation models, there has been a recent shift in training multi-modal foundation models to allow for leveraging information from multiple modalities, either to train a model that simultaneously processes multiple modalities (Mizrahi et al., 2023; Meta, 2024), or to train and bind multiple modality-specific foundation models (Radford et al., 2021; Girdhar et al., 2023; Thapa et al., 2024), particularly with a contrastive objective. Similarly for health applications, there has been a growing interest in cross-modal reconstruction of biosignals (Sarkar & Etemad, 2020), or simultaneously pre-training multiple biosignal modalities (Deldari et al., 2022; Liu et al., 2024; Deldari et al., 2024; Thapa et al., 2024) or wearable metrics (Narayanswamy et al., 2024).
@@
In this work, we adopt ideas from different related work in uni-modal and multi-modal representation learning and knowledge distillation, to distill knowledge from a foundation model for a high-fidelity modality in order to get a foundation model for a low-fidelity modality. Similar to uni-modal foundation models, we train our PPG teacher encoder using masked autoencoding (He et al., 2021) and contrastive learning (Chen et al., 2020), with our own variation of these frameworks. We then transfer knowledge from our PPG teacher encoders to accelerometry student encoders via a cross-modal knowledge distillation framework (Tian et al., 2022) that bears similarity to multi-modal contrastive learning (Radford et al., 2021; Girdhar et al., 2023) and self-distillation for model compression (Fang et al., 2021). There has been prior work on training uni-modal foundation models for PPG and accelerometry (Spathis et al., 2021; Yuan et al., 2023; Abbaspourazad et al., 2024) or multi-modal foundation models for other biosignals (Deldari et al., 2022; Liu et al., 2024; Deldari et al., 2024; Thapa et al., 2024). These multi-modal biosignal foundation models either take multiple modalities as input (Liu et al., 2024; Deldari et al., 2024); or they “bind” multi-modality embeddings into a shared subspace (Deldari et al., 2022; Thapa et al., 2024) similar to (Radford et al., 2021) from randomly-initialized encoders, which is primarily different for our motivation of cross-modal representational knowledge distillation and model compression from an existing high-fidelity teacher encoder. In fact, we show that training both encoders of high-fidelity and low-fidelity modalities together degrades the quality of low-fidelity embeddings (see Section 5.5). In line with our motivation, there has been prior work on leveraging asymmetric information in biosignals, by cross-modal reconstruction of electrocardiogram (ECG) from PPG, for a more accurate estimation of heart rate (Sarkar & Etemad, 2020). We emphasize that our work focuses on accelerometry during low-motion periods where it captures minute cardiovascular signals such as the ballistocardiogram, which is significantly distinct from modeling accelerometry during gross motion for activity recognition, health and fitness (Hallgrímsson et al., 2018; Ni et al., 2019; Spathis et al., 2021). All in all, our work studies unsupervised knowledge distillation across wearable biosignals using large-scale data collected from wearable consumer devices, to develop a foundation model for accelerometry during low-motion periods, that is readily predictive of a wide array of downstream health targets.
Figure 1: Overview of our representational knowledge distillation framework, in two major steps.
Teacher pre-training: we first pre-train an encoder for the high-fidelity biosignal, PPG, using masked autoencoding, this step can be alternated with other self-supervised learning techniques, such as contrastive learning, for which we do an ablation in the text. We turn the multi-channel PPG input into tokens from non-overlapping sequential patches. The encoder Transformer processes only a small proportion of the input tokens (20%), and the decoder Transformer takes the encoder’s output tokens interleaved with learnable mask tokens to reconstruct the PPG signal at masked patches. Cross-modal representational knowledge distillation: we then distill the representational knowledge from the PPG teacher encoder to an accelerometry student encoder using a paired PPG-accelerometry dataset. In each mini-batch of data, we first augment PPG and accelerometry signals independently, the teacher (frozen) and student encoders process these augmented views into embeddings, which are then down-projected to a lower-dimensional subspace. The down-projected embeddings are then aligned using a multi-modal contrastive loss.
@@ 3 Methods and implementation details
Our representational knowledge distillation framework is fully unsupervised and consists of two steps: teacher pre-training and cross-modal representational knowledge distillation as depicted in Figure 1. Below, we go over the details of each of these steps as well as their implementation details.
@@ 3.1 Teacher pre-training
The first step of our representational knowledge distillation framework is to pre-train the PPG teacher encoder via self-supervised learning without any labels. To show that our knowledge distillation framework is agnostic to the pre-training of the PPG teacher encoder and study its efficacy, we investigated two popular pre-training strategies: masked autoencoding (MAE) as the main framework and contrastive learning (CL) as an ablation, which we explain in detail below.
@@
Masked autoencoding: Our masked autoencoding pre-training framework is adopted from the prior work on images (He et al., 2021), for time-series. We turn the multi-channel PPG input (4-channels and 60s-long, as explained in Section 4.1) into patches using non-overlapping fixed-length windows, and then project the patches with a learnable linear tokenizer into tokens, which results in 192 256-D tokens. Sinusoidal positional embeddings are added to the input tokens, then 80% of input tokens are randomly dropped and the 38 kept tokens are passed through the encoder Transformer to get encoder output tokens. Learnable mask tokens, initialized with a 256-D token drawn from a uniform distribution $\sim U(0,1)$, are added back to the encoder output tokens at positions where input tokens where dropped, followed by adding sinusoidal positional embeddings. These 192 new tokens are then processed by the decoder Transformer to generate the decoder output tokens. Finally, with a linear projection, the decoder output tokens are projected to the multi-channel PPG output, with the objective of reconstructing PPG “pixel” values in those indices whose patches/tokens were dropped in the encoder’s input. For maximum learning rate, we used 2e-4 and batch size was set to 512. A complete list of other architectural and training hyperparameters that are shared across methods using Transformers can be found in Appendix Table 4. As a baseline for our distilled accelerometry encoders, we also train masked autoencoders for accelerometry in the same way explained above.
@@
Contrastive learning: While our main pre-training framework for the teacher encoder is masked autoencoding, we perform ablation in regards to the teacher pre-training method and architecture via contrastive learning with EfficientNets. We would like to emphasize that contrastive learning can also be done with Transformer models without any meaningful difference (Appendix Table 13), but we choose EfficientNets for the main results to simultaneously demonstrate an ablation on changing the teacher pre-training strategy and student/teacher architecture. Our contrastive pre-training framework closely follows a prior work for PPG signals (Abbaspourazad et al., 2024), with the major difference that the positive pairs are selected as two augmented views of the same sample. This choice is made to enforce the encoders to contain more segment-level information necessary for the main downstream targets used in this study, as opposed to participant-level positive pair selection which was investigated in the prior work (Appendix Table 14). During pre-training, we augment each sample twice with our stochastic augmentation module that consists of a stochastic cascade of several standard individual augmentations (Iwana & Uchida, 2021) including {cut out: 0.4, magnitude warp: 0.25, add Gaussian noise: 0.25, channel permute: 0.25, time warp: 0.15}, where the values are the assigned probabilities of whether the augmentation is applied for each segment or not. The two augmented views are then passed through a joint-embedding architecture, where one encoder is an exponential moving average ($m=0.99$ for momentum updates) of the other encoder that is updated via backpropagation. The 256-D embeddings of each encoder are then projected to a lower dimensional subspace (128-D) via multi-layer perceptron projection heads. The training objective is maximizing the mutual information of the down-projected embeddings of the two views of the same PPG segment, and minimizing that for different PPG segments, implemented by a regularized InfoNCE loss. We use Kozachenko-Leonenko (KoLeo) differential entropy estimator (Sablayrolles et al., 2019; Oquab et al., 2023) with the weight of $0.1$ for regularization, and temperature of $0.04$ for scaling similarity scores in the InfoNCE loss. For the maximum learning rate, we used 1e-3, while batch size was set to 256. Other common hyperparameters for training EfficientNets is available in Appendix Table 5. As a baseline for our distilled accelerometry encoders, we also train contrastive learned encoders for accelerometry in the same way explained above.

### v2

@@ Abstract
Modern wearable devices can conveniently record various biosignals in the many different environments of daily living, enabling a rich view of individual health. However, not all biosignals are the same: high-fidelity biosignals, such as photoplethysmogram (PPG), contain more physiological information, but require optical sensors with a high power footprint. Alternatively, a lower-fidelity biosignal such as accelerometry has a significantly smaller power footprint and is available in almost any wearable device. While accelerometry is widely used for activity recognition and fitness, it is less explored for health biomarkers and diagnosis. Here, we show that an accelerometry foundation model can predict a wide variety of health targets. To achieve improved performance, we distill representational knowledge from PPG encoders to accelerometery encoders using 20 million minutes of unlabeled data, collected from $\sim$172K participants in the Apple Heart and Movement Study under informed consent. We observe strong cross-modal alignment on unseen data, e.g., $99.2\%$ top-1 accuracy for retrieving PPG embeddings from accelerometry embeddings. We show that distilled accelerometry encoders have significantly more informative representations compared to self-supervised or supervised encoders trained directly on accelerometry data, observed by at least 23%-49% improved performance for predicting heart rate and heart rate variability. We also show that distilled accelerometry encoders are readily predictive of a wide array of downstream health targets, i.e., they are generalist foundation models. We believe accelerometry foundation models for health may unlock new opportunities for developing digital biomarkers from any wearable device.
Machine Learning, ICML
@@ 1 Introduction
The recent growth of wearable devices has empowered individuals to track their health more conveniently and frequently. Wearable devices can record various biosignals and provide individuals with a continuous view of their health metrics in daily environments, a view which is hard to achieve in clinical settings. While this unlocks new exciting opportunities for wearable devices, challenges remain. High-fidelity biosignals, such as photoplethysmogram (PPG) and electrocardiogram (ECG), have substantial physiological information content and are commonly used for developing digital biomarkers. However, they require specific hardware and sensors for recording, therefore are unavailable in some wearable devices. Even when available, it is difficult to collect them frequently due to power constraints of small wearable devices, or reliance on active user engagement for recording (e.g., ECG). On the other hand, a lower-fidelity biosignal, such as accelerometry, is available on almost all wearable devices and has a significantly small power footprint, making it ideal for enabling digital biomarkers efficiently, frequently, and on a variety of wearable devices.
Accelerometry during gross motion is widely used for activity recognition and fitness, where it captures active movements of the body and not direct physiological information. In turn, during low-motion periods, accelerometry is capable of detecting minute physiological movements on the body as a result of heart function, known as ballistocardiogram (Inan et al., 2015; Kim et al., 2016), therefore, it contains complex physiological information. Despite this, accelerometry is less explored for health diagnosis and digital biomarkers on wearable devices compared to other biosignals such as PPG, given that its relationship to diverse health targets is not as well-understood and that it is more susceptible to external artifacts. Examples of how less-structured and noisier accelerometry is compared to PPG can be found in Figure 1. Here, we address this research question by training a generalist foundation model for accelerometry during low-motion, in a large-scale dataset collected under natural living conditions, and demonstrate that accelerometry can predict a diverse set of health targets. See Section 2 for how our work relates to the existing work.
@@
Using unlabeled sensor data collected under informed consent from the large longitudinal Apple Heart and Movement Study (AHMS) (MacRae, 2021; Truslow et al., 2024), we train generalist accelerometry foundation models via cross-modal representational knowledge distillation from PPG. Our contributions are: 1) Accelerometry foundation models for health: We train accelerometry foundation models for health via representational knowledge distillation from PPG using a large-scale dataset with 20M minutes of multi-modal sensor data from $\sim$172K participants. For the first time, we show that a single accelerometry encoder is predictive of a wide array of health targets with a high accuracy. 2) Fully unsupervised representational knowledge distillation framework: While multi-modal modeling and cross-modal reconstruction of biosignals have been done before, we combine and adopt techniques inspired by uni-modal and multi-modal pre-training frameworks from other domains of deep learning to create a fully unsupervised distillation framework for biosignal time-series. Knowledge distillation has been used for improved information extraction from accelerometry, and our end-to-end method being unsupervised is crucial for health applications where labeled data is often limited. 3) Studying representational alignment: We study representational alignment of PPG and accelerometry by doing retrieval analysis. 4) Studying representational information: We study the representational power of accelerometry encoders for critical health targets such as heart rate and heart rate variability, across different available labeled data regimes, as well as for demographic variables and 46 health targets including health conditions, use of medications and lifestyle habits from AHMS survey questions. In addition to these contributions, to support the robustness of the findings, we perform ablation studies to evaluate the efficacy of the encoder architecture (Transformer and EfficientNet), the pre-training strategy of the PPG teacher model (masked autoencoding and contrastive learning), the augmentations, and other training choices.
@@ 2 Related work
Uni-modal and multi-modal representation learning: Unsupervised representation learning techniques, e.g., self-supervised learning, have been proven successful in training generalist models, also known as foundation models, without requiring any explicit labels during training in various domains of deep learning such as natural language processing (Devlin et al., 2019; OpenAI, 2023), computer vision (Chen et al., 2020, 2021; He et al., 2021; Oquab et al., 2023), speech recognition (Baevski et al., 2020, 2022), and health (Cheng et al., 2020; Kostas et al., 2021; Sarkar & Etemad, 2022; Mohsenvand et al., 2020; Gopal et al., 2021; Kiyasseh et al., 2021; Mehari & Strodthoff, 2022; Wu et al., 2020; Spathis et al., 2021; Tang et al., 2021; Yuan et al., 2023; Lai et al., 2023; Abbaspourazad et al., 2024; Liu et al., 2024; Narayanswamy et al., 2024). While most of these works have been primarily on training uni-modal foundation models, there has been a recent shift in training multi-modal foundation models to allow for leveraging information from multiple modalities, either to train a model that simultaneously processes multiple modalities (Mizrahi et al., 2023; Meta, 2024), or to train and bind multiple modality-specific foundation models (Radford et al., 2021; Girdhar et al., 2023; Thapa et al., 2024), particularly with a contrastive objective. Similarly for health applications, there has been a growing interest in cross-modal reconstruction of biosignals (Sarkar & Etemad, 2020), or simultaneously pre-training multiple biosignal modalities (Deldari et al., 2022; Liu et al., 2024; Deldari et al., 2024; Thapa et al., 2024) or their extracted features (Narayanswamy et al., 2024).
@@
In this work, we adopt ideas from different related work in uni-modal and multi-modal representation learning and knowledge distillation, to distill knowledge from a foundation model for a high-fidelity modality in order to get a foundation model for a low-fidelity modality. Similar to uni-modal foundation models, we train our PPG teacher encoder using masked autoencoding (He et al., 2021) and contrastive learning (Chen et al., 2020), with our own variation of these frameworks. We then transfer knowledge from our PPG teacher encoders to accelerometry student encoders via a cross-modal knowledge distillation framework (Tian et al., 2022) that bears similarity to multi-modal contrastive learning (Radford et al., 2021; Girdhar et al., 2023) and self-distillation for model compression (Fang et al., 2021). There has been prior work on training uni-modal foundation models for PPG and accelerometry (Spathis et al., 2021; Yuan et al., 2023; Abbaspourazad et al., 2024; Pillai et al., 2024) or multi-modal foundation models for other biosignals (Deldari et al., 2022; Liu et al., 2024; Deldari et al., 2024; Thapa et al., 2024). These multi-modal biosignal foundation models either take multiple modalities as input (Liu et al., 2024; Deldari et al., 2024); or they “bind” multi-modality embeddings into a shared subspace (Deldari et al., 2022; Thapa et al., 2024) similar to (Radford et al., 2021) from randomly-initialized encoders, which is primarily different for our motivation of cross-modal representational knowledge distillation and model compression from an existing high-fidelity teacher encoder. In fact, we show that training both encoders of high-fidelity and low-fidelity modalities together degrades the quality of low-fidelity embeddings (see Section 5.5). In line with our motivation, there has been prior work on leveraging asymmetric information in biosignals, by cross-modal reconstruction of ECG from PPG, for a more accurate estimation of heart rate (Sarkar & Etemad, 2020). We emphasize that our work focuses on accelerometry during low-motion periods where it captures minute cardiovascular signals such as the ballistocardiogram, which is significantly distinct from modeling accelerometry during gross motion for activity recognition, health and fitness (Hallgrímsson et al., 2018; Ni et al., 2019; Spathis et al., 2021; Xu et al., 2024). Also, self-supervised learning has recently been shown useful for predicting sleep stages from accelerometry (Yuan et al., 2023). All in all, we develop a foundation model for accelerometry during low motion, by distilling embeddings from a pre-trained PPG teacher foundation model, and demonstrate that it is readily predictive of a wide array of downstream health targets.
@@ 3 Methods and implementation details
Our knowledge distillation framework is fully unsupervised and consists of two steps: teacher pre-training and cross-modal representational knowledge distillation as depicted in Figure 1. Below, we go over the details of each of these steps as well as their implementation details.
@@ 3.1 Teacher pre-training
The first step of our representational knowledge distillation framework is to pre-train the PPG teacher encoder via self-supervised learning. To show that our knowledge distillation framework is agnostic to the pre-training of the PPG teacher encoder and study its efficacy, we investigated two popular pre-training strategies: masked autoencoding (MAE) as the main framework and contrastive learning (CL) as an ablation, which we explain in detail below.
@@
Masked autoencoding: Our masked autoencoding pre-training framework is adopted from the prior work on images (He et al., 2021), for time-series. We turn the multi-channel PPG input (4-channels and 60s-long, as explained in Section 4.1) into patches using non-overlapping fixed-length windows, and then project the patches with a learnable linear tokenizer into tokens, which results in 192 256-D tokens. Sinusoidal positional embeddings are added to the input tokens, then 80% of input tokens are randomly dropped and the 38 kept tokens are passed through the encoder Transformer to get encoder output tokens. Learnable mask tokens, initialized with a 256-D token drawn from a uniform distribution $\sim U(0,1)$, are added back to the encoder output tokens at positions where input tokens where dropped, followed by adding sinusoidal positional embeddings. These 192 new tokens are then processed by the decoder Transformer to generate the decoder output tokens. Finally, with a linear projection, the decoder output tokens are projected to the multi-channel PPG output, with the objective of reconstructing PPG “pixel” values in those indices whose patches/tokens were dropped. For maximum learning rate, we used 2e-4 and batch size was set to 512. A complete list of other architectural and training hyperparameters that are shared across methods using Transformers can be found in Appendix Table 4. As a baseline for our distilled accelerometry encoders, we also train masked autoencoders for accelerometry in the same way explained above.
@@
Contrastive learning: While our main pre-training framework for the teacher encoder is masked autoencoding, we perform ablation in regards to the teacher pre-training method and architecture via contrastive learning with EfficientNets. We emphasize that contrastive learning can also be done with Transformer models without any meaningful difference (Appendix Table 13), but we choose EfficientNets for the main results to simultaneously demonstrate an ablation on changing the teacher pre-training strategy and student/teacher architecture. Our contrastive pre-training framework closely follows a prior work for PPG signals (Abbaspourazad et al., 2024). We select positive pairs as two augmented views of the same sample to enforce the encoders to contain more segment-level information necessary for the main downstream targets used in this study (Appendix Table 14). During pre-training, we augment each sample twice with our stochastic augmentation module that consists of a stochastic cascade of several standard individual augmentations (Iwana & Uchida, 2021) including {cut out: 0.4, magnitude warp: 0.25, add Gaussian noise: 0.25, channel permute: 0.25, time warp: 0.15}, where the values are the assigned probabilities of whether the augmentation is applied for each segment or not. The two augmented views are then passed through a joint-embedding architecture, where one encoder is an exponential moving average ($m=0.99$ for momentum updates) of the other encoder that is updated via backpropagation. The 256-D embeddings of each encoder are then projected to a lower dimensional subspace (128-D) via multi-layer perceptron projection heads. The training objective is maximizing the mutual information of the down-projected embeddings of the two views of the same PPG segment, and minimizing that for different PPG segments, implemented by a regularized InfoNCE loss. We use Kozachenko-Leonenko (KoLeo) differential entropy estimator (Sablayrolles et al., 2019; Oquab et al., 2023) with the weight of $0.1$ for regularization, and temperature of $0.04$ for scaling similarity scores in the InfoNCE loss. For the maximum learning rate, we used 1e-3, while batch size was set to 256. Other common hyperparameters for training EfficientNets is available in Appendix Table 5. As a baseline for our distilled accelerometry encoders, we also train contrastive learned encoders for accelerometry in the same way explained above.